# Spectral methods for wedge and corner flows: The Fourier–Kontorovich–Lebedev integral transform

### 11 Mar 2026 v1

@@ III.3.2 Representation of the solution in real space
In the following, we present the methodological steps leading to the system of equations whose solution determines the coefficients $A_{j}$ and $A_{j}^{\dagger}$. Most lengthy expressions are omitted here, as they can be readily obtained from the details provided here, or found in the literature, possibly in slightly different forms depending on the prefactors used in defining these coefficients.
@@ IV.1 Parallel point force
The expressions for $\Delta_{j}$ and $\Delta_{j}^{\dagger}$, for $j\in\{x,y\}$ can be deduced from the original paper by Sano and Hasimoto [60].
@@ IV.2 Transverse point force
$\phi_{x}^{\infty}$ $=-\frac{1}{s}\cos(\beta+\delta)\,$, (44a)
$\phi_{y}^{\infty}$ $=-\frac{1}{s}\sin(\beta+\delta)\,$, (44b)
$\phi_{w}^{\infty}$ $=\frac{1}{s}\cos\delta\,$, (44c)
@@
By requiting vanishing azimuthal velocity on the wedge surface, it follows that
@@
$\Lambda_{w}$ $=-\operatorname{sh}(\beta\nu)\operatorname{sh}(\overline{\alpha}\nu)\operatorname{csh}(\alpha\nu)\cos\delta\,$, (47a)
$\Lambda_{w}^{\dagger}$ $=-\operatorname{ch}(\beta\nu)\operatorname{ch}(\overline{\alpha}\nu)\operatorname{sch}(\alpha\nu)\cos\delta\,.$ (47b)
@@
where $a$ is defined by Eq. (30). These coefficients can be obtained from Sano and Hasimoto [60].
@@ IV.3 Parallel point torque
The expressions for $\Lambda_{j}$, $\operatorname{H}_{j}$, $\Lambda_{j}^{\dagger}$, and $\operatorname{H}_{j}^{\dagger}$, $j\in\{x,y\}$, are obtained by solving the system composed of $\widetilde{\phi}_{r}+\widetilde{\phi}_{r}^{\,\infty}=0$ and Eq. (54) evaluated at $\theta=\pm\alpha$. To evaluate the FKL transform of $\phi_{w}/r$, we use the properties listed in Tab. 1. These coefficients can be found in Ref. 70.
@@ IV.4 Transverse point torque
$\phi_{r}^{\infty}$ $=\frac{z}{s^{3}}\,\sin(\theta-\beta-\delta)\,$, (56a)
$\phi_{\theta}^{\infty}$ $=\frac{z}{s^{3}}\,\cos(\theta-\beta-\delta)\,$, (56b)
$\phi_{z}^{\infty}$ $=-\frac{1}{s^{3}}\left(r\sin(\theta-\beta-\delta)+\rho\sin\theta\right),$ (56c)
$\phi_{w}^{\infty}$ $=\frac{z}{s^{3}}\,\sin\delta\,.$ (56d)
We note that by construction
@@
$\widetilde{\phi}_{r}^{\infty}$ $=-\rho\sin(\theta-\beta-\delta)\,\mathscr{T}_{i\nu}\left\{\frac{\partial}{\partial z}\frac{1}{s}\right\},$ (58a)
$\widetilde{\phi}_{\theta}^{\infty}$ $=-\rho\cos(\theta-\beta-\delta)\,\mathscr{T}_{i\nu}\left\{\frac{\partial}{\partial z}\frac{1}{s}\right\},$ (58b)
$\widetilde{\phi}_{z}^{\infty}$ $=\left(\sin\delta\,\rho\,\frac{\partial}{\partial\rho}-\cos\delta\operatorname{sgn}(\theta-\beta)\,\frac{\partial}{\partial a}\right)\mathscr{T}_{i\nu}\left\{\frac{1}{s}\right\},$ (58c)
$\widetilde{\phi}_{w}^{\infty}$ $=-\rho\sin\delta\,\mathscr{T}_{i\nu}\left\{\frac{\partial}{\partial z}\frac{1}{s}\right\}.$ (58d)
@@
$\Delta_{w}$ $=-\operatorname{sh}(\beta\nu)\operatorname{sh}\left(\overline{\alpha}\nu\right)\operatorname{csh}(\alpha\nu)\sin\delta\,$, (61a)
$\Delta_{w}^{\dagger}$ $=-\operatorname{ch}(\beta\nu)\operatorname{ch}\left(\overline{\alpha}\nu\right)\operatorname{sch}(\alpha\nu)\sin\delta\,.$ (61b)
@@
$\Lambda_{z}$ $=\nu\operatorname{ch}(\beta\nu)\operatorname{sh}\left(\overline{\alpha}\nu\right)\operatorname{csh}(\alpha\nu)\cos\delta\,$, (62a)
$\Lambda_{z}^{\dagger}$ $=\nu\operatorname{sh}(\beta\nu)\operatorname{ch}\left(\overline{\alpha}\nu\right)\operatorname{sch}(\alpha\nu)\cos\delta\,.$ (62b)
@@
The expressions for $\Delta_{j}$ and $\Delta_{j}^{\dagger}$, $j\in\{x,y\}$, can be obtained by solving the system of equations formed by $\widetilde{\phi}_{r}+\widetilde{\phi}_{r}^{\,\infty}=0$ together with Eq. (60), both evaluated at $\theta=\pm\alpha$. We use the properties listed in Tab. 1 to evaluate the FKL transforms of $(z/r)\,\phi_{z}$ and $\phi_{w}/r$. These coefficients can be found in Ref. 70.
@@ V Discussion and concluding remarks
We presented an overview of the application of the FKL transform to address problems in fluid mechanics related to the fundamental solutions of the Stokes equations in wedge-shaped geometries. We detailed the analytical approach based on the Papkovich–Neuber representation of the solution for incompressible flow, involving four harmonic functions. The approach combines the use of a Fourier transform along the axial direction and a KL transform along the radial direction. Upon applying the FKL transform, the problem reduces to ordinary differential equations in the polar angle. The solution is expressed as the sum of the free-space contribution and a complementary solution required to satisfy the no-slip boundary conditions of vanishing velocity at the wedge surfaces. We showed that, depending on the orientation of the point force or point torque, the solution takes a particular structure. The solution in real space is obtained by inverse transform, where the integration with respect to the axial wavenumber is carried out analytically, leaving a single improper integral in the radial wavenumber $\nu$.
@@
The present review focuses primarily on the mathematical formulation of the problems under investigation, beginning with the relevant properties and functional relationships of the FKL transform. The fundamental building block for solving Stokes flow problems in wedge geometries due to singularities is the FKL transform of $1/s$, whose partial derivatives can be used to construct higher-order singularities such as a rotlet. This framework can also be applied to problems related to the dynamics of active microswimmers through the evaluation of the force dipole, source dipole, force quadrupole, and related singularities [81, 82, 83, 84, 85, 86]. The results presented here correspond to the case of no-slip boundary conditions imposed on both walls. The same methodology can be extended to examine free-slip or stress-free boundary conditions, as well as mixed combinations of different types of boundary conditions.
@@
Since the kernel is a modified Bessel function of the second kind, exhibiting asymptotically exponential decay for large indices; c.f. Eq. (21), numerical integration converges rapidly for large radial wavenumbers. Accordingly, numerical integration is typically performed with the upper integration limit truncated at 100, ensuring that the no-slip boundary conditions are satisfied while contributions from larger wavenumbers remain negligible. Increasing the upper integration limit further does not change the results, making this choice an efficient balance between accuracy and computational cost. In the planar boundary limit, corresponding to a wedge opening angle of $\pi$, the integrals admit a closed analytical form, recovering the previously known solutions for a point force [87] and point torque [88].
@@
The methodology presented here can, in principle, be extended to related geometries with angular or corner-like boundaries. For example, it could be adapted to study flows in finite wedges with bounded radial or axial extent. This can be achieved using dual integral approach formulations, which the author has previously applied to study the effects of finite-sized boundaries on various Stokes flows [89, 90, 91, 92, 93]. The main requirement is that the geometry allows a separation of variables compatible with the transform approach, so that the flow can be expressed in terms of harmonic functions. While the transforms and boundary conditions may become more involved in these cases, the framework remains applicable and provides a systematic way to construct solutions for singularity-driven flows in more complex confined geometries.

### 09 Apr 2026 v2

@@ III.3.2 Representation of the solution in real space
In the following, we present the methodological steps leading to the system of equations whose solution determines the coefficients $A_{j}$ and $A_{j}^{\dagger}$. Some lengthy expressions associated with the solution for a point torque are omitted here, as they can be readily obtained from the details provided here, or found in Ref. 71.
In the limit $r\to 0$, relevant for describing the flow behavior near the apex of the wedge, the kernel function given by Eq. (37) assumes the asymptotic form, cf. Appendix
$\hskip-7.11317pt\mathcal{K}_{i\nu}(r,z)\simeq\frac{2\operatorname{csh}(\pi\nu)}{\sqrt{\pi}}\operatorname{Im}\left\{\frac{\Gamma\left(\frac{1}{2}-i\nu\right)}{\Gamma\left(1-i\nu\right)}\frac{\left(\rho^{2}+z^{2}\right)^{i\nu-\frac{1}{2}}}{\left(\rho r\right)^{i\nu}}\right\}$ (39)
with $\Gamma$ denoting the Euler gamma function.
The asymptotic expression of $\phi_{j}$ as $r\to 0$ can be obtained by performing the integration over $\nu$ using the method of residues. The integration can be performed in the upper half of the complex plane using a contour composed of the real axis and a large counterclockwise semicircular arc of radius $R$, with $R$ taken to infinity. In some cases, a small clockwise contour around the origin must be included when zero is a pole. Since there are infinitely many poles, the leading-order behavior is captured by considering only the first few. Not all poles can be determined analytically, which prevents a complete asymptotic solution for an arbitrary wedge opening angle.
@@ IV.1 Parallel point force
The expressions for the coefficients $\Delta_{j}$ and $\Delta_{j}^{\dagger}$, for $j\in\{x,y\}$ are obtained as
$\Delta_{x}=h_{1}\sin\alpha\operatorname{ch}(\alpha\nu)\,,\quad\Delta_{y}=h_{2}\cos\alpha\operatorname{ch}(\alpha\nu)\,$, (45)
together with
$\Delta_{x}^{\dagger}=-h_{2}\sin\alpha\operatorname{sh}(\alpha\nu)\,,\quad\Delta_{y}^{\dagger}=-h_{1}\cos\alpha\operatorname{sh}(\alpha\nu)\,$, (46)
where $h_{1}$ and $h_{2}$ are given by
$h_{1}$ $=4\Gamma_{+}\operatorname{sh}(\pi\nu)\big(\cos\alpha\sin\beta\operatorname{sh}(\alpha\nu)\operatorname{ch}(\beta\nu)$
$\quad-\sin\alpha\cos\beta\operatorname{ch}(\alpha\nu)\operatorname{sh}(\beta\nu)\big)\,$, (47a)
$h_{2}$ $=4\Gamma_{-}\operatorname{sh}(\pi\nu)\big(\sin\alpha\cos\beta\operatorname{sh}(\alpha\nu)\operatorname{ch}(\beta\nu)$
$\quad-\cos\alpha\sin\beta\operatorname{ch}(\alpha\nu)\operatorname{sh}(\beta\nu)\big)\,$, (47b)
where
$\hskip-2.84544pt\Gamma_{\pm}^{-1}=\left(\operatorname{sh}(2\alpha\nu)\pm\nu\sin(2\alpha)\right)\left(\operatorname{ch}(2\alpha\nu)\mp\cos(2\alpha)\right).$ (48)
@@ IV.2 Transverse point force
$\phi_{x}^{\infty}$ $=-\frac{1}{s}\cos(\beta+\sigma)\,$, (49a)
$\phi_{y}^{\infty}$ $=-\frac{1}{s}\sin(\beta+\sigma)\,$, (49b)
$\phi_{w}^{\infty}$ $=\frac{1}{s}\cos\sigma\,$, (49c)
@@
By requiring vanishing azimuthal velocity on the wedge surface, it follows that
@@
$\Lambda_{w}$ $=-\operatorname{sh}(\beta\nu)\operatorname{sh}(\overline{\alpha}\nu)\operatorname{csh}(\alpha\nu)\cos\sigma\,$, (52a)
$\Lambda_{w}^{\dagger}$ $=-\operatorname{ch}(\beta\nu)\operatorname{ch}(\overline{\alpha}\nu)\operatorname{sch}(\alpha\nu)\cos\sigma\,.$ (52b)
@@
where $a$ is defined by Eq. (30). Defining the abbreviations $\zeta_{\pm}=\alpha\pm\beta$, $\xi_{\pm}=\alpha\pm\beta\pm\sigma$, $\beta_{\pm}=\pi\pm\beta$, and $\tau=\beta+\sigma$, the final results can be represented in the form
$\Lambda_{x}$ $=\Pi_{+}\left(b\cos\tau+\nu\left({a}_{-}+{c}_{-}\operatorname{ch}(\alpha\nu)\right)\sin\alpha\right),$ (54a)
$\Lambda_{y}$ $=\Pi_{-}\left(b\sin\tau+\nu\left({a}_{+}-{c}_{+}\operatorname{ch}(\alpha\nu)\right)\cos\alpha\right),$ (54b)
$\Lambda_{x}^{\dagger}$ $=\Pi_{-}\left(b^{\dagger}\cos\tau-\nu\left({a}_{+}^{\dagger}-{c}_{+}\operatorname{sh}(\alpha\nu)\right)\sin\alpha\right),$ (54c)
$\Lambda_{y}^{\dagger}$ $=\Pi_{+}\left(b^{\dagger}\sin\tau-\nu\left({a}_{-}^{\dagger}+{c}_{-}\operatorname{sh}(\alpha\nu)\right)\cos\alpha\right).$ (54d)
The coefficients associated with $\rho\,\partial/\partial\rho$ in Eq. (32) can be cast in the form
$\hskip-5.69046pt\operatorname{H}_{x}$ $=-\Pi_{+}d_{-}\sin\alpha\operatorname{ch}(\alpha\nu)\,,\,\operatorname{H}_{y}=\Pi_{-}d_{+}\cos\alpha\operatorname{ch}(\alpha\nu)\,$, (55a)
$\hskip-5.69046pt\operatorname{H}_{x}^{\dagger}$ $=-\Pi_{-}d_{+}\sin\alpha\operatorname{sh}(\alpha\nu)\,,\,\operatorname{H}_{y}^{\dagger}=\Pi_{+}d_{-}\cos\alpha\operatorname{sh}(\alpha\nu)\,$, (55b)
where
$\Pi_{\pm}^{-1}=\operatorname{sh}(2\alpha\nu)\pm\nu\sin(2\alpha)\,.$ (56)
We have introduced the following expressions to present the results more concisely
${a}_{\pm}$ $=\cos\xi_{+}\operatorname{sh}\left(\beta_{-}\nu\right)\pm\cos\xi_{-}\operatorname{sh}\left(\beta_{+}\nu\right),$ (57a)
${a}_{\pm}^{\dagger}$ $=\cos\xi_{+}\operatorname{ch}\left(\beta_{-}\nu\right)\pm\cos\xi_{-}\operatorname{ch}\left(\beta_{+}\nu\right),$ (57b)
together with
${c}_{\pm}$ $=Z_{\pm}\left(\cos\zeta_{-}\operatorname{ch}\left(\zeta_{+}\nu\right)\pm\cos\zeta_{+}\operatorname{ch}\left(\zeta_{-}\nu\right)\right),$ (58a)
$d_{\pm}$ $=Z_{\pm}\left(\sin\zeta_{-}\,\operatorname{sh}\left(\zeta_{+}\nu\right)\pm\sin\zeta_{+}\,\operatorname{sh}\left(\zeta_{-}\nu\right)\right),$ (58b)
where
$Z_{\pm}=\frac{2\cos\sigma\operatorname{sh}(\pi\nu)}{\operatorname{ch}\left(2\alpha\nu\right)\pm\cos(2\alpha)}\,.$ (59)
Moreover,
$b$ $=2\operatorname{ch}(\alpha\nu)\operatorname{sh}(\overline{\alpha}\nu)\operatorname{sh}(\beta\nu)\,$, (60a)
$b^{\dagger}$ $=2\operatorname{sh}(\alpha\nu)\operatorname{ch}(\overline{\alpha}\nu)\operatorname{ch}(\beta\nu)\,$, (60b)
where again $\overline{\alpha}=\pi-\alpha\,.$
@@ IV.3 Parallel point torque
The expressions for $\Lambda_{j}$, $\operatorname{H}_{j}$, $\Lambda_{j}^{\dagger}$, and $\operatorname{H}_{j}^{\dagger}$, $j\in\{x,y\}$, are obtained by solving the system composed of $\widetilde{\phi}_{r}+\widetilde{\phi}_{r}^{\,\infty}=0$ and Eq. (66) evaluated at $\theta=\pm\alpha$. To evaluate the FKL transform of $\phi_{w}/r$, we use the properties listed in Tab. 1. These coefficients are provided in Eq. (32) of Ref. 71.
@@ IV.4 Transverse point torque
$\phi_{r}^{\infty}$ $=\frac{z}{s^{3}}\,\sin(\theta-\beta-\sigma)\,$, (68a)
$\phi_{\theta}^{\infty}$ $=\frac{z}{s^{3}}\,\cos(\theta-\beta-\sigma)\,$, (68b)
$\phi_{z}^{\infty}$ $=-\frac{1}{s^{3}}\left(r\sin(\theta-\beta-\sigma)+\rho\sin\theta\right),$ (68c)
$\phi_{w}^{\infty}$ $=\frac{z}{s^{3}}\,\sin\sigma\,.$ (68d)
By construction
@@
$\widetilde{\phi}_{r}^{\infty}$ $=-\rho\sin(\theta-\beta-\sigma)\,\mathscr{T}_{i\nu}\left\{\frac{\partial}{\partial z}\frac{1}{s}\right\},$ (70a)
$\widetilde{\phi}_{\theta}^{\infty}$ $=-\rho\cos(\theta-\beta-\sigma)\,\mathscr{T}_{i\nu}\left\{\frac{\partial}{\partial z}\frac{1}{s}\right\},$ (70b)
$\widetilde{\phi}_{z}^{\infty}$ $=\left(\sin\sigma\,\rho\,\frac{\partial}{\partial\rho}-\cos\sigma\operatorname{sgn}(\theta-\beta)\,\frac{\partial}{\partial a}\right)\mathscr{T}_{i\nu}\left\{\frac{1}{s}\right\},$ (70c)
$\widetilde{\phi}_{w}^{\infty}$ $=-\rho\sin\sigma\,\mathscr{T}_{i\nu}\left\{\frac{\partial}{\partial z}\frac{1}{s}\right\}.$ (70d)
@@
$\Delta_{w}$ $=-\operatorname{sh}(\beta\nu)\operatorname{sh}\left(\overline{\alpha}\nu\right)\operatorname{csh}(\alpha\nu)\sin\sigma\,$, (73a)
$\Delta_{w}^{\dagger}$ $=-\operatorname{ch}(\beta\nu)\operatorname{ch}\left(\overline{\alpha}\nu\right)\operatorname{sch}(\alpha\nu)\sin\sigma\,.$ (73b)
@@
$\Lambda_{z}$ $=\nu\operatorname{ch}(\beta\nu)\operatorname{sh}\left(\overline{\alpha}\nu\right)\operatorname{csh}(\alpha\nu)\cos\sigma\,$, (74a)
$\Lambda_{z}^{\dagger}$ $=\nu\operatorname{sh}(\beta\nu)\operatorname{ch}\left(\overline{\alpha}\nu\right)\operatorname{sch}(\alpha\nu)\cos\sigma\,.$ (74b)
@@
The expressions for $\Delta_{j}$ and $\Delta_{j}^{\dagger}$, $j\in\{x,y\}$, can be obtained by solving the system of equations formed by $\widetilde{\phi}_{r}+\widetilde{\phi}_{r}^{\,\infty}=0$ together with Eq. (72), both evaluated at $\theta=\pm\alpha$. We use the properties listed in Tab. 1 to evaluate the FKL transforms of $(z/r)\,\phi_{z}$ and $\phi_{w}/r$. These coefficients are provided in Eq. (42) of Ref. 71.
@@ V Discussion and concluding remarks
Figure 2: Scaled magnitude of the fluid velocity on the upper and lower walls induced by an axial point force, obtained via numerical integration with the upper limit truncated at $\nu_{\text{max}}$. The lines represent exponential fits. The velocity is scaled by $F/(8\pi\eta\rho)$. Here, $\alpha=\pi/6$, $\beta=\alpha/2$, and $r/\rho=z/\rho=1/2$.
@@
We presented an overview of the application of the FKL transform to address problems in fluid mechanics related to the fundamental solutions of the Stokes equations in wedge-shaped geometries. We detailed the analytical approach based on the Papkovich–Neuber representation of the solution for incompressible flow, involving four harmonic functions. The approach combines the use of a Fourier transform along the axial direction and a KL transform along the radial direction. Upon applying the FKL transform, the problem reduces to ordinary differential equations in the polar angle. The solution is expressed as the sum of the free-space contribution and a complementary solution required to satisfy the no-slip boundary conditions of vanishing velocity at the wedge surfaces. We showed that, depending on the orientation of the point force or point torque, the solution assumes a particular structure. The solution in real space is obtained by inverse transform, where the integration with respect to the axial wavenumber is carried out analytically, leaving a single improper integral in the radial wavenumber $\nu$.
@@
The present review focuses primarily on the mathematical formulation of the problems under investigation, beginning with the relevant properties and functional relationships of the FKL transform. The fundamental building block for solving Stokes flow problems in wedge geometries due to singularities is the FKL transform of $1/s$, whose partial derivatives can be used to construct higher-order singularities such as a rotlet. This framework can also be applied to problems related to the dynamics of active microswimmers through the evaluation of the force dipole, source dipole, force quadrupole, and related singularities [81, 82, 83, 84, 85, 86]. The results presented here correspond to the case of no-slip boundary conditions imposed on both walls. The same method can be extended to examine free-slip or stress-free boundary conditions, as well as mixed combinations of different types of boundary conditions.
@@
Since the kernel is a modified Bessel function of the second kind, which exhibits asymptotically exponential decay for large indices; c.f. Eq. (21), the numerical integration converges rapidly for large radial wavenumbers. Accordingly, the numerical integration is typically performed with the upper integration limit truncated at a specified maximum radial wavenumber, ensuring that the no-slip boundary conditions are satisfied while contributions from larger wavenumbers remain negligible. Figure 2 illustrates this exponential decay for a parallel point force using representative values for the wedge geometry. The boundary condition on the lower wall at $\theta=-\alpha$ is satisfied with a smaller truncation of the upper limit, owing to the faster exponential decay. Similar overall dynamics are observed for a point force aligned along the transverse direction or for a point torque. Typically, setting $\nu_{\text{max}}=100$ ensures an error below $10^{-10}$, making it suitable for evaluating the improper integrals. Increasing the upper integration limit further does not change the results, making this choice an efficient balance between accuracy and computational cost. In the planar boundary limit, corresponding to a wedge opening angle of $\pi$, the integrals admit a closed analytical form, thereby recovering the previously known solutions for a point force [87] and a point torque [88].
@@
The FKL transform is primarily a technical method for addressing flows near corners, providing an efficient framework for solving such problems. Unlike the method of successive reflections, which is a classical singularity-based approach relying on repeated image systems and often resulting in cumbersome series, the FKL approach expresses the solution in terms of rapidly converging integrals. This not only simplifies the computational implementation but also allows the effects of wedge boundaries to be incorporated directly into the kernel, offering a clear and compact representation of the flow induced by point singularities. However, the method is subject to certain limitations. The approach is specifically tailored to wedge-shaped domains and relies on the separability of the geometry in cylindrical coordinates; extending it to more complex or irregular geometries would require substantial modifications. Additionally, the method is formulated for isolated point singularities; extending it to fully distributed forces, such as those around particles, would require additional treatments. The method presented here can, in principle, be extended to related geometries with corner-like boundaries. For example, it could be adapted to study flows in finite wedges with bounded radial or axial extent. This can be achieved using dual integral equation formulations, which the author has previously applied to study the effects of finite-sized boundaries on various Stokes flows [89, 90, 91, 92, 93]. The main requirement is that the geometry allows a separation of variables compatible with the transform approach, so that the flow can be expressed in terms of harmonic functions. While the transforms and boundary conditions may become more involved in these cases, the framework remains applicable and provides a systematic way to construct solutions for singularity-driven flows in more complex confined geometries.
The solution outlined in the present work has the only approximation stemming from the numerical evaluation of the integral over $\nu$, which can be made arbitrarily accurate by choosing a sufficiently large upper limit and appropriate quadrature. To our knowledge, no other analytical or semi-analytical method exists for wedge-shaped geometries, as the combination of a point singularity and wedge boundaries makes closed-form solutions extremely challenging. In contrast, numerical approaches such as boundary element or finite-volume methods require discretization of the fluid domain or boundaries, involving two- or three-dimensional integrations, which are computationally more demanding. By reducing the problem to one-dimensional integrals over the radial wavenumber, our method achieves a significant gain in efficiency while maintaining high accuracy, which is particularly advantageous when evaluating multiple configurations or performing parameter studies in wedge geometries.
Considerable efforts have been devoted in the literature to solving the Stokes equations for fluid flows and the Navier–Cauchy equations for linear elasticity using boundary element methods. These methods rely on singular solutions to construct numerical solutions by discretizing the boundaries of the domain. In wedge-like geometries, this generally requires meshing both the surfaces of the walls and the objects within the domain, which can be computationally intensive. By directly incorporating the Green’s functions outlined in this work, which satisfy the no-slip boundary conditions, into the boundary element formulation, the numerical implementation is greatly simplified. In this approach, the wedge walls do not need to be discretized, leading to a substantial reduction in computational cost. The solution detailed here can therefore be used to simulate particle dynamics in wedge-like geometries, as well as the behavior of inclusions in elastic media, providing a versatile tool for modeling in these complex geometries.